# Exploration Behavior of Untrained Policies

### v1

@@ 2.2 Smooth Networks Yield Ballistic Trajectories
Let $\pi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a stochastic policy sampled from a neural network prior, such that the realized sample function is $L_{\pi}$-Lipschitz. Let the deterministic transition dynamics $s_{t+1}=f(s_{t},a_{t})$ satisfy Assumption 1, and also be Lipschitz continuous with respect to the action space: $|f(s,a)-f(s,a^{\prime})|\leq L_{f}|a-a^{\prime}|$ and state space, $|f(s,a)-f(s^{\prime},a)|\leq L_{s}|s-s^{\prime}|$ for all $s,s^{\prime}\in\mathcal{S}$ and $a,a^{\prime}\in\mathcal{A}$. Let the constant $c=\pi_{\theta}(s_{0})$ denote the initial action and let $L_{s}=1$ (the $L_{s}\neq 1$ regime is dominated by exponential time dependence, and is further discussed in the Appendix). Then for any initial state $s_{0}$ the agent’s trajectory approximately follows the straight-line path $ct$, up to a bounded error:
@@
Although the previous lemma holds for all times, a linear approximation holds only for short timescales, $t\ll\frac{2c}{\delta L_{a}L_{\pi}}$, beyond which, deviations caused by curvature (resulting from non-linearities in the deep neural net) become large enough to dominate the effective mean velocity, $c$.
The Lipschitz constant of a neural policy can be computed [Virmaux and Scaman(2018), Latorre et al.(2020)Latorre, Rolland, and Cevher, Bhowmick et al.(2021)Bhowmick, D’Souza, and Raghavan] or roughly bounded a priori to make Lemma 1 implementable. Although the assumptions and structure of the above dynamics may be limiting in some environments, they provide a starting point to develop intuition for the relationship between architecture, dynamics, and “depth” [^1] In the sense of “deep exploration” [Osband et al.(2016)Osband, Blundell, Pritzel, and Van Roy]. of trajectories. We will focus on generalizing these results in future work.
@@
Figure 1: Though trajectories can change direction (as evidence in the plot on the right), on short timescales, the trajectories can be well-approximated with linear drift.
Figure 2: Standard MLPs used for deep RL (with two hidden layers, $256$ hidden nodes, ReLU activation) produce ballistic trajectories.
@@ 2.3 Random Policy Networks as Gaussian Processes
With this setup, we can leverage the following closed-form expression for the policy in the infinite-width limit (cf. also Eq. 1 of [Meronen et al.(2021)Meronen, Trapp, and Solin] and surrounding discussion for more details):
@@
Note that these properties are complementary to that of the fixed policy, whose architecture is typically Lipschitz, implying the ballistic trajectories discussed above.
@@ 2.4 Case Study: ReLU Network and the Induced State Distribution
where $\cos\theta=\frac{\langle s,s^{\prime}\rangle}{|s||s^{\prime}|}$. This kernel is not stationary (i.e. it does not only depend on a radial distance between inputs), and induces a diffusion term that grows with $|s|^{2}$. Thus, far from the origin, exploration becomes more diffuse. Stationary kernels (e.g. Radial Basis Functions [Buhmann(2000)]) depending on $|s-s^{\prime}|$ on the other hand would result in constant diffusion coefficient, which may be of interest in some environments.
In the special case of the one hidden layer, infinitely wide ReLU, we take weights and biases drawn from centered Gaussians with variance $\sigma_{w},\sigma_{b}$, respectively. In this case, the policy is mean zero and has diffusion coefficient $K(s,s)=\Sigma(s)=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{\pi}\|s\|^{2}$. For the simplest linear dynamics $s_{t+1}=s_{t}+\pi(s_{t})$, we can model the stochastic policy as a random walk in state space. In the continuous time (equivalently “small action”) limit, the Fokker-Planck equation can be employed to study the dynamical distribution $p(s,t)$ over state space:
@@
With radial symmetry $p_{\infty}(s)=f(r)$, $r=\|s\|$, we obtain the solution:
@@
This describes a heavy-tailed stationary distribution (resembling Cauchy or power-law distributions) and is normalizable if $\sigma_{b}>0$. Note that as $\sigma_{b}^{2}\to 0$, the tails become increasingly heavy, and exploration is more likely to reach distant regions of state space.

### v3

@@ 2.2 Smooth Networks Yield Ballistic Trajectories
Let $\pi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a stochastic policy sampled from a neural network prior, such that the realized sample function is $L_{\pi}$-Lipschitz. Let the deterministic transition dynamics $s_{t+1}=f(s_{t},a_{t})$ satisfy Assumption 1, and also be Lipschitz continuous with respect to the action space: $|f(s,a)-f(s,a^{\prime})|\leq L_{a}|a-a^{\prime}|$ and state space, $|f(s,a)-f(s^{\prime},a)|\leq L_{s}|s-s^{\prime}|$ for all $s,s^{\prime}\in\mathcal{S}$ and $a,a^{\prime}\in\mathcal{A}$.
Let the constant $c=\pi_{\theta}(s_{0})$ denote the initial action and let $L_{s}=1$ (the $L_{s}\neq 1$ regime is dominated by exponential time dependence, and is further discussed in the Appendix). Then for any initial state $s_{0}\in\mathcal{S}$, the agent’s trajectory approximately follows the straight-line path $ct$, up to a bounded error:
@@
Although the previous lemma holds for all times, a linear approximation to the trajectory: ${s_{t}=s_{0}+ct+\mathcal{O}(t^{2})}$ holds only for short timescales, $t\ll\frac{2c}{\delta L_{a}L_{\pi}}$, beyond which, deviations caused by curvature (resulting from non-linearities in the deep neural net) become large enough to dominate the effective mean velocity, $c$.
The Lipschitz constant of a neural policy can be computed [Virmaux and Scaman(2018), Latorre et al.(2020)Latorre, Rolland, and Cevher, Bhowmick et al.(2021)Bhowmick, D’Souza, and Raghavan] or roughly bounded a priori to make Lemma 1 practically implementable. Although the assumptions and structure of the above dynamics may be limiting for some environments, the result provides a starting point to develop an intuition for the relationship between architecture, dynamics, and “depth” [^1] In the sense of “deep exploration” [Osband et al.(2016)Osband, Blundell, Pritzel, and Van Roy]. of trajectories. We will focus on generalizing these results in future work.
@@
Figure 1: Though trajectories can change direction (as observed in the plot on the right), on short timescales, the trajectories can be well-approximated with a dominant linear drift.
Figure 2: Standard MLPs used for deep RL (with two hidden layers, $256$ hidden nodes, ReLU activation) produce “ballistic” trajectories.
@@ 2.3 Random Policy Networks as Gaussian Processes
With this setup, we can leverage the following closed-form expression for the policy in the infinite-width limit (cf. also Eq. 1 of [Meronen et al.(2021)Meronen, Trapp, and Solin] and surrounding discussion therein for more details):
@@
Note that these properties are complementary to that of the fixed policy, whose architecture is typically Lipschitz, implying the ballistic trajectories discussed in Section 2.2.
@@ 2.4 Case Study: ReLU Network and the Induced State Distribution
where $\cos\theta=\frac{\langle s,s^{\prime}\rangle}{|s||s^{\prime}|}$. This kernel is not stationary (i.e. it does not only depend on a radial distance between inputs), and induces a diffusion term that grows with $|s|^{2}$. Thus, far from the origin, action selection becomes less correlated. Stationary kernels (e.g. Radial Basis Functions [Buhmann(2000)]) depending on $|s-s^{\prime}|$ on the other hand would result in a constant diffusion coefficient, which may be of interest in some environments.
In the special case of one hidden layer, infinitely wide, with ReLU non-linearities, we take weights and biases drawn from centered Gaussians with variance $\sigma_{w},\sigma_{b}$, respectively. In this case, the policy is mean zero and has diffusion coefficient $K(s,s)=\Sigma(s)=\sigma_{b}^{2}+\frac{\sigma_{w}^{2}}{\pi}\|s\|^{2}$. For the simplest linear dynamics $s_{t+1}=s_{t}+\pi(s_{t})$, we can model the stochastic policy as a random walk in state space. In the continuous time (equivalently “small action”) limit, the Fokker-Planck equation can be employed to study the dynamical distribution $p(s,t)$ over state space:
@@
With radial symmetry $p_{\infty}(s)=f(r)$, $r=\|s\|$, we obtain the following solution:
@@
This describes a heavy-tailed stationary distribution (resembling Cauchy or power-law distributions) and is normalizable if $\sigma_{b}>0$. Note that as $\sigma_{b}^{2}\to 0$, the tails become increasingly heavy, and the agent is more likely to reach distant regions of state space.
@@ 3 Discussion
Acknowledgements
JA acknowledges funding support from the PDT Partners Machine Learning Conference Grant and the NSF through Award No. PHY-2425180. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).